### v2

# Data-driven $h^{2}$ model reduction for linear discrete-time systems

Abstract
We present a new framework of $h^{2}$ optimal model reduction for linear discrete-time systems. Our main contribution is to create optimal reduced-order models in the $h^{2}$-norm sense directly from the measurement data alone, without using the information of the original system. In particular, we focus on the fact that the gradient of the $h^{2}$ model reduction problem is expressed using the discrete-time Lyapunov equation and the discrete-time Sylvester equation, and derive the data-driven gradient. In the proposed algorithm, the initial point can be chosen as the output of the existing data-driven methods. In numerical experiments, we demonstrate that, for a modeling task in neuroscience, our method constructs a reduced-order model that outperforms DMDc in terms of the $h^{2}$-norm.
Index Terms:
Data-driven model order reduction, Discrete-time dynamical systems
I Introduction
Model Order Reduction (MOR) is a powerful tool for compressing high-dimensional dynamical systems, enabling faster simulation and control without sacrificing essential dynamics. In systems and control theory, MOR generates reduced-order models (ROMs) that are capable of reproducing the input-output behavior of large-scale dynamical systems with high accuracy. The resulting ROMs have the advantage of being easier to manipulate and control, unlike the inherently larger systems. There are MOR methods based on Singular Value Decomposition (SVD) [1, 2, 3, 4, 5], Krylov subspaces (or moment matching methods) [6, 4, 7, 8], and the optimization of the $h^{2}$ (or $H^{2}$) norm [7, 9, 10, 11, 12]. These classical model reduction methods based on the state-space description of the system are known as model-based MOR. On the other hand, when a state-space description of the system is not available or computational modeling is difficult, it is desirable to create ROMs using the measurement data alone.
Data-driven model reduction constructs the ROMs that are capable of reproducing the input-output behavior of large-scale dynamical systems with high accuracy directly from the measurement data alone. Unlike traditional approaches, an explicit parametric system-identification step is not required; therefore, no additional identification-model error is introduced. Several well-known data-driven MOR methods have been proposed due to the increasing use of data. The data-driven balanced truncation method [13, 14, 15, 16] is one of the common data-driven model reduction inspired by the balanced truncation method, which is a model-based MOR. In this method, balanced truncation is performed by using the measurement data. For example, [15] performs MOR by estimating the Gramians from the data information. Another well-known approach is data-driven interpolatory model reduction. This approach includes the Loewner framework [17, 18, 19, 20, 21, 22, 23] and the AAA algorithm [24, 25]. Other data-driven methods, such as proper orthogonal decomposition (POD) [26], dynamic mode decomposition (DMD) [27, 28], or operator inference [29, 30] can also be employed. These approaches are the leading data-driven model reduction, but they may not yield optimal ROMs in the $h^{2}$-norm sense. In such cases, the output of the resulting reduced system may not approximate that of the original system.
In this paper, we propose a data-driven model reduction method for linear discrete-time systems. We focus on obtaining the optimal ROMs in the $h^{2}$-norm sense, using the measurement time-domain data. The contributions of this research are the following two points:
1) We propose a new framework for data-driven model reduction for linear discrete-time systems. We focus on obtaining the ROMs that perform well in terms of the $h^{2}$-norm, directly from the measurement data. In our approach, there is no need to use system identification methods, as in traditional approaches. Furthermore, under certain assumptions about the measurement data, we derive the gradients of the optimization problem characterized by the data.
2) We develop an algorithm for the proposed method and conduct numerical experiments. The proposed algorithm can be hybridized with existing data-driven MOR. In numerical experiments, we demonstrate that, for a modeling task in neuroscience, our method constructs a ROM that outperforms DMDc in terms of the $h^{2}$-norm.
The remainder of the paper is structured as follows. Section II describes the model-based $h^{2}$ MOR problem for the discrete-time systems. We describe the problem setting of this study in Section III. Section IV describes the data-driven $h^{2}$ MOR method proposed in this paper and its algorithm is explained in Section V. Section VI presents the results of numerical experiments. We conclude and discuss future work in Section VII.
Notation: We denote the imaginary unit by $\mathrm{i}$. We also denote the Frobenius norm, transpose, trace, the Moore-Penrose pseudo-inverse, the $i$-th eigenvalue, the $i$-th row vector of the matrix $A$, and the $(i,j)$-th element of the matrix $A$ by $\|A\|_{F}$, $A^{\top}$, $\operatorname{tr}A$, $A^{{\dagger}}$, $\lambda_{i}(A)$, and $A(i,:)$, $[A]_{i,j}$, respectively. In addition to that, we denote the complex conjugate transpose of $A\in\mathbb{C}^{n\times n}$ by $A^{*}\in\mathbb{C}^{n\times n}$.
II Preliminaries
In this section, we summarize a model-based $h^{2}$ MOR [4, 7, 9, 10] for linear time-invariant (LTI) discrete-time dynamical systems
[…]
with transfer function $H(z):=C(zI_{n}-A)^{-1}B$, where $k$ indicates the temporal iteration and $u_{k}\in\mathbb{R}^{m}$, $x_{k}\in\mathbb{R}^{n}$, and $y_{k}\in\mathbb{R}^{p}$ denote the input, state, and output at time $t_{k}$, respectively. The matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, and $C\in\mathbb{R}^{p\times n}$ are constant matrices, which are known.
In this study, we assume that $C=I_{n}$. This means $p=n$. That is, we consider a system in which all state data are observed as outputs. This assumption is commonly made in the context of DMD, as explained in Section III. In addition, we assume that the system $\Sigma$ is asymptotically stable, i.e. all eigenvalues of $A\in\mathbb{R}^{n\times n}$ lie inside the unit circle. Such a matrix $A$ is called a stable matrix.
The ROM of system (1) is defined as
[…]
with transfer function $\hat{H}(z):=\hat{C}(zI_{r}-\hat{A})^{-1}\hat{B}$, where $\hat{x}_{k}\in\mathbb{R}^{r}$, $\hat{y}_{k}\in\mathbb{R}^{n}$, and $r\ll n$. In this context, $\hat{C}$ can be regarded as a matrix that reconstructs the dimensionally reduced state data back into the original output space, and it is employed to compare original system (1) and ROM (2).
II-A Model-Based $h^{2}$ Model Order Reduction
To track the output error $x_{k}-\hat{C}\hat{x}_{k}$ between (1) and (2), we focus on finding a ROM (2) that minimizes the approximation error $H-\hat{H}$ in the $h^{2}$-norm under the constraints that $(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r}$ and $\hat{A}$ is stable. Note that the $h^{2}$-norm of the discrete-time system $\Sigma$ is defined as
[…]
The $h^{2}$ MOR problem for the original system (1) is written as
$&\mathop{\rm minimize}\limits&&\|H-\hat{H}\|_{h^{2}}\\ &\mathop{\rm subject\ to}\limits&&(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r},\\ &&&\hat{A}\>\text{is stable}.\end{aligned}$ (3)
Proposition 1 ([4][10]).
Consider the asymptotically stable systems (1) with $C=I_{n}$ and (2). Then, the objective function of (3) can be rewritten as
$\|H-\hat{H}\|_{h^{2}}^{2}$ $=\operatorname{tr}(\Sigma_{c})+f(\hat{A},\hat{B},\hat{C})$
$=\operatorname{tr}(B^{\top}\Sigma_{o}B)+f(\hat{A},\hat{B},\hat{C}).$
Here,
$f(\hat{A},\hat{B},\hat{C})$ $:=\operatorname{tr}(\hat{C}P\hat{C}^{\top})-2\operatorname{tr}(R\hat{C}^{\top})$
$=\operatorname{tr}(\hat{B}^{\top}Q\hat{B})+2\operatorname{tr}(B^{\top}S\hat{B}),$
and $\Sigma_{c}$, $\Sigma_{o}$, $P$, $Q$, $R$, and $S$ are the unique solutions of the following equations:
$A\Sigma_{c}A^{\top}+BB^{\top}$ $=\Sigma_{c},$ (4)
$A^{\top}\Sigma_{o}A+I_{n}$ $=\Sigma_{o},$ (5)
$\hat{A}P\hat{A}^{\top}+\hat{B}\hat{B}^{\top}$ $=P,$ (6)
$\hat{A}^{\top}Q\hat{A}+\hat{C}^{\top}\hat{C}$ $=Q,$ (7)
$AR\hat{A}^{\top}+B\hat{B}^{\top}$ $=R,$ (8)
$A^{\top}S\hat{A}-\hat{C}$ $=S.$ (9)
where (6) and (7) are the discrete-time Lyapunov equations and (8) and (9) are the discrete-time Sylvester equations.
From Proposition 1, (3) can be rewritten as
$&\mathop{\rm minimize}\limits&&f(\hat{A},\hat{B},\hat{C})\\ &\mathop{\rm subject\ to}\limits&&(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r},\\ &&&\hat{A}\>\text{is stable}.\end{aligned}$ (10)
A modification as in (10) eliminates the need to compute the large-Lyapunov equations as in (4) and (5).
II-B Gradients of (10)
Let us derive the gradients of $f$ versus $\hat{A}$, $\hat{B}$, and $\hat{C}$. We define a gradient as follows.
Definition 1 ([9]).
The gradients of a real scalar smooth function $f(X)$ of a real matrix variable $X\in\mathbb{R}^{n\times p}$ is the real matrix $\nabla_{X}f(X)\in\mathbb{R}^{n\times p}$ defined by
$[\nabla_{X}f(X)]_{i,j}=\frac{\partial}{\partial X_{i,j}}f(X),i=1,\ldots,n,j=1,\ldots,p.$
Proposition 2 ([9][10]).
The gradients $\nabla_{\hat{A}}f$, $\nabla_{\hat{B}}f$, and $\nabla_{\hat{C}}f$ of $f$ are given by
$\nabla_{\hat{A}}f(\hat{A},\hat{B},\hat{C})=2(Q\hat{A}P+S^{\top}AR),$ (11)
$\nabla_{\hat{B}}f(\hat{A},\hat{B},\hat{C})=2(S^{\top}B+Q\hat{B}),$ (12)
$\nabla_{\hat{C}}f(\hat{A},\hat{B},\hat{C})=2(\hat{C}P-R),$ (13)
where $P$, $Q$, $R$, and $S$ are the solutions to (6), (7), (8), and (9), respectively.
When the gradients $(\nabla_{\hat{A}}f,\nabla_{\hat{B}}f,\nabla_{\hat{C}}f)$ are all zero, the stationary conditions derived in [10], i.e., Wilson’s first order necessary conditions, are satisfied. In this study, we derive these gradients directly from the measurement data without using the constant matrices $A$ and $B$ of the system (1).
III Problem formulation
We formulate the problem addressed in this study under the assumption that the matrices $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ for the original system (1) are unknown. Note that this is a different setting from Section II. Instead, assume that $N$ measurement data sets obtained from the true system (1) is given by
$\begin{pmatrix}x_{1,1}&x_{1,2}&\dots&x_{1,L}\\ x_{2,1}&x_{2,2}&\dots&x_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N,1}&x_{N,2}&\dots&x_{N,L}\end{pmatrix},\quad\begin{pmatrix}u_{1,1}&u_{1,2}&\dots&u_{1,L}\\ u_{2,1}&u_{2,2}&\dots&u_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N,1}&u_{N,2}&\dots&u_{N,L}\end{pmatrix}$ (14)
where $L$ is a total number of measurement instances for the data sets $i=1,2,\ldots,N$. For (14), $x_{i,k+1}$, $x_{i,k}$, and $u_{i,k}$ satisfy
$x_{i,k+1}=Ax_{i,k}+Bu_{i,k},$ (15)
where $i=1,\ldots,N$ and $k=1,\ldots,L-1$.
Remark 1.
While the assumption that state data is directly observable is restrictive, this is identical to the standard DMD framework [27], which has been applied in various fields such as neuroscience [31], fluid dynamics [32], epidemiology [33], and financial engineering [34].
In this paper, we address the following problem.
Problem 1.
Develop an algorithm for solving the $h^{2}$ MOR problem described in (10) in a data-driven manner. Specifically, construct a stable ROM using the measurement data (14), under the condition that the matrices $A$ and $B$ in (1) are unknown.
Note that the gradients as expressed in Proposition 2 are not directly applicable due to the unknown nature of matrices $A$ and $B$.
Problem 1 is important for the following reason: we examine a system with transfer function $H_{\textrm{SI}}$, which is obtained by a system identification method. Then,
$\|H-\hat{H}\|_{h^{2}}$ $=\|H-H_{\textrm{SI}}+H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$
$\leq\|H-H_{\textrm{SI}}\|_{h^{2}}+\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}.$ (16)
In traditional approaches, the minimization of the error $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ is considered. However, if errors $\|H-H_{\textrm{SI}}\|_{h^{2}}$ arise from system identification, minimizing $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ may still not result in satisfactory ROMs in the $h^{2}$-norm. Furthermore, if the system identified is unstable, $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ cannot be defined and the $h^{2}$ MOR may not be performed. To overcome these problems, this paper proposes a MOR method to optimize the left-hand side of (III) directly from the measurement data.
IV Data-driven $h^{2}$ Model Reduction for Discrete-time Systems
In this section, we describe a data-driven $h^{2}$ MOR for (10). By making some assumptions on the measurement data of the original system (1), we show that ROMs can be obtained directly from the data when the matrices $A$ and $B$ of (1) are unknown.
IV-A Data-driven Reconstruction of the State Equation
To solve the discrete-time Sylvester equation (8) in a data-driven manner, we represent the system’s state equation (15) using the measurement data (14). First, we collect the data in the matrices
$X:=$ $\big{(}x_{1,1},x_{1,2},\ldots,x_{1,L-1},x_{2,1},x_{2,2},\ldots,x_{2,L-1},$ (17)
$\ldots,x_{N,1},x_{N,2},\ldots,x_{N,L-1}\big{)}^{\top}\in\mathbb{R}^{(NL-N)\times n},$
$X_{\mathrm{p}}:=$ $\big{(}x_{1,2},x_{1,3},\ldots,x_{1,L},x_{2,2},x_{2,3},\ldots,x_{2,L},$
$\ldots,x_{N,2},x_{N,3},\ldots,x_{N,L}\big{)}^{\top}\in\mathbb{R}^{(NL-N)\times n},$
$U:=$ $\big{(}u_{1,1},u_{1,2},\ldots,u_{1,L-1},u_{2,1},u_{2,2},\ldots,u_{2,L-1},$
$\ldots,u_{N,1},u_{N,2},\ldots,u_{N,L-1}\big{)}^{\top}\in\mathbb{R}^{(NL-N)\times m},$
where the $i$-th row of $X_{\mathrm{p}}$ corresponds to the $i$-th row of $X$, shifted by one step according to the system’s state equation.
Next, we represent the system’s state equation using the measurement data. Using (15) and the data matrices (17),
$X_{\mathrm{p}}=XA^{\top}+UB^{\top}.$ (18)
Multiplying this equation by $X^{\top}$ from the right yields
$X_{\mathrm{p}}X^{\top}=XA^{\top}X^{\top}+UB^{\top}X^{\top}.$ (19)
Here (19) can be rewritten as
$X_{\mathrm{p}}X^{\top}$ $=XZ_{A}^{\top}+UZ_{B}^{\top},$ (20)
where
$Z_{A}$ $:=XA,$ (21)
$Z_{B}$ $:=XB.$ (22)
Furthermore, transposing (19) yields
$XU_{B}^{\top}$ $=XX_{\mathrm{p}}^{\top}-Z_{A}X^{\top},$ (23)
where
$U_{B}$ $:=UB^{\top}.$ (24)
Note that $Z_{A}$ corresponds to the matrix of state data for the dual system in (15).
In general, the solutions $Z_{A}$, $Z_{B}$, and $U_{B}$ of (20) and (23) are not uniquely determined as in (21), (22), and (24). We denote the solutions obtained by solving (20) and (23) as $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$, respectively.
To prove the uniqueness, we consider the following assumption.
Assumption.
Conditions for the Proposed Method:
  1. (A1)
    The spectra of $X^{\dagger}Z_{A,\text{data}}$ and $\hat{A}^{-1}$ are disjoint.
  2. (A2)
    The spectra of $X^{\dagger}(X_{\mathrm{p}}-U_{B,\text{data}})$ and $\hat{A}^{-1}$ are disjoint.
  3. (A3)
    The eigenvalues of $\hat{A}$ are inside the unit circle and strictly positive, i.e, $\forall i\in\{1,2,\ldots,r\},0<|\lambda_{i}(\hat{A})|<1$.
  4. (A4)
    $\operatorname{rank}\begin{bmatrix}X&U\end{bmatrix}=n+m$.
  5. (A5)
    $\operatorname{rank}X=n$.
  6. (A6)
    $\operatorname{rank}U=m$.
Remark 2.
Assumptions (A1)-(A6) may be easily achieved for the following reasons.
Assumptions (A1)-(A3) are conditions related to the eigenvalues of the $\hat{A}$ matrix. As described in Section V, the proposed algorithm allows the $\hat{A}$ matrix to be constructed to satisfy these assumptions. For details, refer to Section V.
Assumptions (A4)-(A6) are specifically related to the measurement data. If the number of data sets $N$ and measurement instances $L$ satisfy $NL-N\geq n+m$, and the vectors $(x_{i,k},u_{i,k})^{\top}$, $x_{i,k}$, and $u_{i,k}$ are linearly independent for each $i=1,\ldots,N$ and $k=1,\ldots,L$, respectively, then the assumptions hold. Furthermore, these assumptions are related to the persistency of excitation condition of the input [35]. Specifically, when the input $u$ possesses sufficient persistent excitation, the Hankel matrix of $U$ becomes full rank, and the rank of $X$ coincides with the state dimension $n$. Therefore, for example, as shown in [36], if an input $u$ that satisfies the persistency of excitation condition is generated, both $X$ and $U$ are guaranteed to be full rank.
IV-B Construction of Gradient-like Matrices from Measurement Data
In this subsection, we show that the gradient-like matrices corresponding to Proposition 2 can be derived from the data. To obtain the gradient-like matrices, we solve the discrete-time Sylvester equations (8) and (9) in a data-driven manner, following [37][38]. Multiplying (8) and (9) by the data matrix $X$ from the left, from (18), (21), (22), and (24), we obtain
$XAR\hat{A}^{\top}+XB\hat{B}^{\top}=XR$
$\Leftrightarrow Z_{A}R\hat{A}^{\top}+Z_{B}\hat{B}^{\top}=XR,$ (25)
$XA^{\top}S\hat{A}-X\hat{C}=XS$
$\Leftrightarrow(X_{\mathrm{p}}-U_{B})S\hat{A}-X\hat{C}=XS.$ (26)
Furthermore, multiplying $X^{\dagger}$ from the left-hand side of (25) and (26), they can be rewritten as
$X^{\dagger}Z_{A}R\hat{A}^{\top}+X^{\dagger}Z_{B}\hat{B}^{\top}=X^{\dagger}XR,$ (27)
$X^{\dagger}(X_{\mathrm{p}}-U_{B})S\hat{A}-X^{\dagger}X\hat{C}=X^{\dagger}XS,$ (28)
where the solutions $R$ and $S$ of (27) and (28) are denoted by $R_{\text{data}}$ and $S_{\text{data}}$, respectively.
Lemma 1.
Let $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$, where $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$ are the solutions of (21), (22), and (24), respectively. Suppose that Assumptions (A1)-(A3), and (A5) hold. Then, the solutions $R$ and $S$ to (27) and (28) exist uniquely.
Proof.
It is shown using the well-known unique solution existence theorem for the Sylvester equation [39]. See Appendix. $\Box$
Using the relation $AR=(R-B\hat{B}^{\top})(\hat{A}^{\dagger})^{\top}$ from (8), and the solutions $R_{\text{data}}$ of (27) and $S_{\text{data}}$ of (28), the gradient-like matrices for (10) are calculated as
~$\tilde{\nabla}_{\hat{A}}f(\hat{A},\hat{B},\hat{C}):=2(Q\hat{A}P+(S_{\text{data}}^{\top}R_{\text{data}}-S_{B,\text{data}}^{\top}\hat{B}^{\top})(\hat{A}^{\dagger})^{\top}),$ (29)
~$\tilde{\nabla}_{\hat{B}}f(\hat{A},\hat{B},\hat{C}):=2(S_{B,\text{data}}^{\top}+Q\hat{B}),$ (30)
~$\tilde{\nabla}_{\hat{C}}f(\hat{A},\hat{B},\hat{C}):=2(\hat{C}P-R_{\text{data}}),$ (31)
where $S_{B,\text{data}}$ denotes the solution $S_{B}$ to the equation
$US_{B}=U_{B}S_{\text{data}},$ (32)
and, as stated in Lemma 3, it coincides with $B^{\top}S_{\text{data}}$ under certain conditions. In general, however, (29)-(31) do not coincide with (11)-(13) because $R_{\text{data}}$ and $S_{\text{data}}$ are different from the solutions $R$ and $S$ in (8) and (9). Using such inaccurate gradients could lead to convergence to incorrect solutions. In what follows, we derive sufficient conditions on the data for the gradient-like matrices to coincide with the gradients of Proposition 2.
IV-C Data-driven $h^{2}$ Model Reduction
This subsection describes the data-driven model reduction for the case where $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown. To solve (10) numerically, we compute the gradients of Proposition 2 in a data-driven manner.
Lemma 2.
Let $NL-N\geq n+m$. Suppose that Assumptions (A4)-(A6) hold. Then, (20) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$.
Proof.
We use the condition for the existence of a unique solution to the linear equation. See Appendix. $\Box$
Lemma 2 means that under assumptions, $Z_{A}$ given by (21) and $Z_{B}$ given by (22) can be expressed using input and state data of system (1). In other words, $Z_{A}$ and $Z_{B}$ can be expressed as $Z_{A,\text{data}}$ and $Z_{B,\text{data}}$ without using the unknown matrices $A$ and $B$.
Lemma 3.
Let $NL-N\geq n+m$. Suppose that Assumptions (A4)-(A6) hold. Then, (23) and (32) have the unique solutions (24) and
$S_{B}=B^{\top}S_{\text{data}},$ (33)
respectively.
Proof.
It is proved following the same approach as in Lemma 2. See Appendix. $\Box$
The following theorem presents the conditions under which the gradient-like matrices, obtained through a data-driven approach, become the gradients for (10).
Theorem 1.
Let $NL-N\geq n+m$. Suppose that Assumptions (A1)-(A6) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (10) specified in Proposition 2, that is, ~$\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold.
Proof.
It follows from Lemmas 1, 2, and 3. See Appendix. $\Box$
Theorem 1 states that the gradients for (10) are computed based on the data given by (17) under Assumptions (A1)-(A6). In fact, $R_{\text{data}}$, $S_{\text{data}}$, and $S_{B,\text{data}}$ in (29), (30), and (31) can be calculated by solving (27), (28), and (32). Moreover, $(Z_{A},Z_{B})$ and $U_{B}$ in (27), (28), and (32) can be calculated by solving (20) and (23). As mentioned earlier, if Assumptions (A1)-(A6) are not satisfied, it is possible that the gradient obtained from the data differs from the true gradients as shown in Proposition 2. When Assumptions (A1)-(A6) regarding the data are satisfied and the gradients from Theorem 1 are all zero, the stationary point condition described in Section II is satisfied. This fact becomes crucial in the construction of the proposed algorithm, as discussed in Section V.
Remark 3.
Our framework assumes that the full state is directly measured, that is, $p=n$ and, without loss of generality, $C=I_{n}$. Under this setting we have $y_{i,k}=x_{i,k}$, and therefore the state sequence is exactly known and Theorem 1 provides the true gradients of (10). If only a part of the state is observed ($p<n$) or, more generally, if $C\neq I_{n}$, the state $x_{i,k}$ must be reconstructed, which inevitably introduces approximation errors. In such a partially observed setting the matrices ~$\tilde{\nabla}_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f$ no longer coincide with the true gradients. Future work will address this case.
IV-D Gradients of (10) for the case that $A$ is unknown and $B$ is known
We derive the gradients for the case where the matrix $B$ for (1) is known. Unlike the setting in Subsection IV-C, the conditions for the gradient-like matrices to coincide with the gradients of Proposition 2 are different.
Corollary 1.
Let $NL-N\geq n$. Suppose that Assumptions (A1)-(A3), and (A5) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (10) specified in Proposition 2, that is, ~$\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold true.
Proof.
The proof follows directly from that of Theorem 1. $\Box$
The next corollary is the condition under which the gradients of Theorem 1 and Corollary 1 coincide.
Corollary 2.
Under Assumptions (A1)-(A6), the gradients (29)-(31) with unknown $B\in\mathbb{R}^{n\times m}$ coincide with the gradients with known $B\in\mathbb{R}^{n\times m}$.
Proof.
From Lemma 3, we obtain $\nabla_{\hat{A}}f=2(Q\hat{A}P+(S^{\top}R-S_{B}^{\top}\hat{B}^{\top})(\hat{A}^{-1})^{\top})=2(Q\hat{A}P+S^{\top}(R-B\hat{B}^{\top})(\hat{A}^{-1})^{\top})$, $\nabla_{\hat{B}}f=2(S_{B}^{\top}+Q\hat{B})=2(S^{\top}B+Q\hat{B})$. $\Box$
V Gradient-based data-driven $h^{2}$ MOR algorithm
In this section, we present Algorithm 1 which solves the problem (10) when $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown.
We briefly describe Algorithm 1. First, initial reduced matrices $\hat{\theta}_{1}=(\hat{A}_{(1)},\hat{B}_{(1)},\hat{C}_{(1)})$ are created to satisfy Assumption (A3). In step 1 and 2 of Algorithm 1, $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$, which are needed to solve the discrete-time Sylvester equations (27) and (28) with $(Z_{A},Z_{B},U_{B})=(Z_{A,\text{data}},Z_{B,\text{data}},U_{B,\text{data}})$, are generated. At each iteration, after computing the gradients, the reduced matrices are updated so that the objective function for (10) becomes smaller. In the while statement of Algorithm 1, the backtracking method is executed.
Here, we show that for a sufficiently small step size $\alpha$, the updated matrix $\bar{A}$ satisfies Assumption 3. Since the set of stable matrices $\mathcal{\hat{D}}_{\text{stab}}:=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},|\lambda_{i}(\hat{A})|<1\}$ is open [40], and the set of nonsingular matrices $\mathcal{\hat{D}}_{\text{nonsing}}:=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},|\lambda_{i}(\hat{A})|>0\}$ is also open due to the continuity of the determinant, their intersection $\mathcal{\hat{D}}:=\mathcal{\hat{D}}_{\text{stab}}\bigcap\mathcal{\hat{D}}_{\text{nonsing}}=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},0<|\lambda_{i}(\hat{A})|<1\}$ is open by the basic topological axiom that the finite intersection of open sets is open. Hence $\mathcal{\hat{D}}$, which denotes the set of matrices satisfying Assumption 3, is an open subset of $\mathbb{R}^{r\times r}$. Thus, there exists $\varepsilon(\hat{A}_{(\ell)})>0$ such that $\left\{\bar{A}\in\mathbb{R}^{r\times r}\,\middle|\,\|\bar{A}-\hat{A}_{(\ell)}\|_{F}<\varepsilon(\hat{A}_{(\ell)})\right\}\subset\hat{\mathcal{D}}$. Hence for the gradient $\nabla_{\hat{A}}f(\hat{A}_{(\ell)})$, we choose
$0<\alpha_{\ell}<\alpha^{\star}_{\ell}\quad\text{with}\quad\alpha^{\star}_{\ell}:=\frac{\varepsilon(\hat{A}_{(\ell)})}{\|\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\|_{F}}.$
Then $\|\hat{A}_{(\ell)}-\alpha_{\ell}\nabla_{\hat{A}}f(\hat{A}_{(\ell)})-\hat{A}_{(\ell)}\|_{F}=\alpha_{\ell}\|\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\|_{F}<\varepsilon(\hat{A}_{(\ell)})$ and therefore $\bar{A}:=\hat{A}_{(\ell)}-\alpha_{\ell}\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\in\hat{\mathcal{D}}.$
In addition, since $c>0$ and $D>0$, there exists $\alpha$ satisfying the Armijo condition. Furthermore, from a numerical perspective, the spectrum of $\hat{A}^{-1}$ does not intersect with that of $X^{\dagger}Z_{A,\text{data}}$ and $X^{\dagger}(X_{\mathrm{p}}-U_{B,\text{data}})$. That is, $\alpha$ can be determined by the backtracking method such that the Armijo condition and Assumptions (A1)- (A3) are satisfied.
Remark 4.
Since (10) is a non-convex optimization problem, the choice of the initial point is important. As the initial point for Algorithm 1, it is possible to use ROMs obtained through the following existing data-driven model reduction methods.
  • DMDc [27][28]
  • Loewner framework [17, 22, 21, 20, 19]
  • Data-driven balanced truncation [15, 16, 14]
Using ROMs obtained from these methods as initial points for Algorithm 1 often enables the construction of high-quality ROMs in the sense of the $h^{2}$ norm. However, these data-driven methods require different types of data, and therefore, they may not be directly applicable in some cases.
The convergence of Algorithm 1 is guaranteed by the following proposition.
Proposition 3.
Suppose that the sequence $\hat{\theta}_{\ell}:=(\hat{A}_{(\ell)},\hat{B}_{(\ell)},\hat{C}_{(\ell)})$ generated by Algorithm 1 is bounded. In addition, we assume that each step-size $\alpha_{\ell}$ is chosen by the backtracking method with parameters $10<c_{1},\rho<$ 1 in Algorithm 1. Then the whole sequence $\{\hat{\theta}_{\ell}\}$ converges to a stationary point of problem (10).
Proof.
The backtracking method gives $f(\hat{\theta}_{\ell+1})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\cdot\|\nabla f(\hat{\theta}_{\ell})\|^{2}_{F}$, therefore $\{f(\hat{\theta}_{\ell})\}$ is monotone and bounded. Summation yields $\sum_{\ell=1}^{\infty}\alpha_{\ell}\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<\infty$.
Since $f$ is continuously differentiable and $10<\rho<$ 1, there exists the lower bound $\alpha_{\ell}\geq\underline{\alpha}>0$[41]. Hence $\underline{\alpha}\cdot\sum_{\ell=1}^{\infty}\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<\infty$ and therefore $\|\nabla f(\hat{\theta}_{\ell})\|_{F}\to 0$.
Because $\|\hat{\theta}_{\ell+1}-\hat{\theta}_{\ell}\|_{F}=\alpha_{\ell}\|\nabla f(\hat{\theta}_{\ell})\|_{F}$, $\{\hat{\theta}_{\ell}\}$ is Cauchy sequence and converges. Continuity of $\nabla f$ implies that the limit point $\hat{\theta}^{\star}$ satisfies $\nabla f(\hat{\theta}^{\star})=0$, i.e. it is stationary. $\Box$
Lastly, we provide the computational complexity analysis of Algorithm 1.
Theorem 2.
We assume that the number of trials of the backtracking method and the number of iterations of Algorithm 1 are $\mathcal{O}(k)$ and $\mathcal{O}(l)$, respectively. Then, the overall computational complexity of Algorithm 1 is given by
$\mathcal{O}((NL-N)^{2}(n+m)+((NL-N)n^{2}+n^{3}+r^{3})kl).$
Proof.
See Appendix. $\Box$
A widely–used alternative for constructing a ROM directly from snapshot data is DMDc [28, 27]. Given the snapshot matrix $X\in\mathbb{R}^{(NL-N)\times n}$, $X_{\text{p}}\in\mathbb{R}^{(NL-N)\times n}$, $U\in\mathbb{R}^{(NL-N)\times m}$ by (17), the computational complexity by DMDc [28, 27] is $\mathcal{O}\!\bigl{(}(NL-N)^{2}(r+m)+(NL-N)nr\bigr{)}$. Although Theorem 2 indicates that the proposed method incurs a higher computational complexity than DMDc, it can generate a stable ROM with a smaller $h^{2}$ error, as demonstrated in Section VI.
Algorithm 1 Solve (10) for the case that $A$ and $B$ are unknown
0:  Measurement data $(x_{i,k}$, $u_{i,k})$, $(i=1,2,\ldots,N,k=1,2,\ldots,L)$, dimension $r$ for (2), initial step-size $\alpha$, Armijo parameter $c_{1}$, search control parameter $\rho$, tolerance $tol$, the initial real matrices $\hat{\theta}_{1}=(\hat{A}_{(1)},\hat{B}_{(1)},\hat{C}_{(1)})$ which satisfy Assumption (A3)
0:  $\hat{\theta}=(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r}$
1:  Generate $Z_{A,\text{data}}$ and $Z_{B,\text{data}}$ from (20) with $(Z_{A},Z_{B})=(Z_{A,\text{data}},Z_{B,\text{data}})$
2:  Generate $U_{B,\text{data}}$ from (23) with $(Z_{A},U_{B})=(Z_{A,\text{data}},U_{B,\text{data}})$
3:  for $\ell=1,2,\ldots$ do
4:     Solve (6) and (7) for $P$ and $Q$ with $\hat{\theta}=\hat{\theta}_{\ell}$
5:     Solve (27) and (28) for $R$ and $S$ with $\hat{\theta}=\hat{\theta}_{\ell}$
6:     Calculate the gradients $\nabla_{\hat{A}}f(\hat{\theta}_{\ell})$, $\nabla_{\hat{B}}f(\hat{\theta}_{\ell})$, and $\nabla_{\hat{C}}f(\hat{\theta}_{\ell})$ with $S_{B}$ obtained by (32)
7:     Define $\nabla f(\hat{\theta}_{\ell}):=-(\nabla_{\hat{A}}f(\hat{\theta}_{\ell}),\nabla_{\hat{B}}f(\hat{\theta}_{\ell}),\nabla_{\hat{C}}f(\hat{\theta}_{\ell}))$
8:     if $\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<tol$ then break
9:     $\alpha_{\ell}=\alpha$
10:     while 1 do
11:        $\bar{\theta}:=\hat{\theta}_{\ell}+\alpha_{\ell}d$
12:        Solve (6) for $P$ with $\hat{\theta}=\bar{\theta}$
13:        Solve (27) for $R$ with $\hat{\theta}=\bar{\theta}$
14:        if $f(\bar{\theta})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\cdot\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}$ and $\forall i,0<|\lambda_{i}(\bar{A})|<1$ then $\hat{\theta}_{\ell+1}:=\bar{\theta}$ break
15:        $\alpha_{\ell}\leftarrow\rho\alpha_{\ell}$
16:     end while
17:  end for
VI Application to Computational Neuroscience
In Computational Neuroscience, it is crucial to elucidate how the neuronal network of Caenorhabditis elegans (C. elegans) regulates behavioral states [31, 42, 43]. Specifically, the behavioral states of C. elegans can be analyzed by observing the fluorescence intensity of neuronal groups [42, 43]. Moreover, it is known that changes in oxygen concentration affect the neural activity of C. elegans [42, 43].
Therefore, in this experiment, as in [31], we model the relationship between the neural activity of C. elegans, regarded as the state, and changes in oxygen concentration, regarded as the control input, using a state-space model. Furthermore, we demonstrate that the ROMs obtained by the proposed data-driven method can reconstruct the original system more accurately in the sense of the $h^{2}$-norm compared to DMDc [28, 27].
VI-A Problem setting
We consider the following system with noise $\epsilon_{k}$:
[…]
where $k=1,\ldots,L-1$.
First, we describe the measurement data for the original system (34). The data utilized in this study are accessible at https://osf.io/a64uz/ and are linked to two prior experimental papers [43, 31]. Using the data corresponding to [43], we constructed the state data $X$ and $X_{\mathrm{p}}$ with noise and $\tilde{X}$ and $\tilde{X}_{\mathrm{p}}$ without noise, as well as the control input data $U$, as defined by (17) and (34). More specifically, we utilized the calcium imaging data from the npr-1 mutant strain, which is sensitive to changes in oxygen concentration, during the pre-lethargus stage of this strain as the state data. The changes in oxygen concentration were used as the control input. Note that, in this dataset, the parameters for (17) were set as $N=1$, $L=4045$, $n=114$, and $m=1$. Using the MATLAB randn command, the noise $\epsilon_{k}$ was generated. Before generating noise, we fixed the seed with rng(0,’twister’) to ensure exact reproducibility.
To incorporate broader temporal information, we consider a time delay of $h_{x}=9$ steps for the state data $X$ and $h_{u}=2$ steps for the input data $U$. Accordingly, we construct the following Hankel matrices $X^{(h_{x})}$ from $X$ in (17) with $\bar{L}:=L-h_{x}$:
$X^{(h_{x})}=\begin{bmatrix}x_{1}&x_{2}&\cdots&x_{\bar{L}}\\ x_{2}&x_{3}&\cdots&x_{\bar{L}+1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{h_{x}}&x_{h_{x}+1}&\cdots&x_{\bar{L}+h_{x}-1}\end{bmatrix}\in\mathbb{R}^{h_{x}n\times\bar{L}}.$
In the similar way, $\tilde{X}^{(h_{x})}\in\mathbb{R}^{h_{x}n\times\bar{L}}$, $X^{(h_{x})}_{\mathrm{p}}\in\mathbb{R}^{h_{x}n\times\bar{L}}$, $\tilde{X}^{(h_{x})}_{\mathrm{p}}\in\mathbb{R}^{h_{x}n\times\bar{L}}$, and $U^{(h_{u})}\in\mathbb{R}^{h_{u}m\times\bar{L}}$ are constructed from $X$, $X_{\mathrm{p}}$, and $U$ in (17), respectively. Therefore, we consider the following extended state-space model for (34):
[…]
where $x_{k}^{(h_{x})}$, $\tilde{x}_{k}^{(h_{x})}$, and $u_{k}^{(h_{u})}$ represent the $k$-th columns of the matrices $X^{(h_{x})}$, $\tilde{X}^{(h_{x})}$, and $U^{(h_{u})}$, respectively, and $\epsilon_{k}^{(h_{x})}=(\epsilon_{k},\dots,\epsilon_{k+h_{x}-1})$. Using (35), $A^{(h_{x})}$ and $B^{(h_{u})}$ matrices are calculated from $X^{(h_{x})}$, $X^{(h_{x})}_{\mathrm{p}}$, and $U^{(h_{u})}$ as follows:
$A^{(h_{x})}$ $=\left[\tilde{X}^{(h_{x})}_{\mathrm{p}}\begin{bmatrix}\tilde{X}^{(h_{x})}\\ U^{(h_{u})}\end{bmatrix}^{\dagger}\right](:,1:h_{x}n),$
$B^{(h_{u})}$ $=\left[\tilde{X}^{(h_{x})}_{\mathrm{p}}\begin{bmatrix}\tilde{X}^{(h_{x})}\\ U^{(h_{u})}\end{bmatrix}^{\dagger}\right](:,h_{x}n+1:h_{x}n+h_{u}m).$
We then describe other parameters. The initial step-size $\alpha$, Armijo parameter $c_{1}$, and search control parameter $\rho$ required for the back tracking method were $\alpha=1$, $c=10^{-4}$, and $\rho=0.5$, respectively. The tolerance $tol$ was $tol=10^{-3}$.
In this experiment, we demonstrate that the ROMs obtained by the proposed method approximates the true system (35) more accurately in terms of frequency response compared to the ROMs obtained by DMDc. Specifically, the objective function is the relative error $\frac{\|H-\hat{H}\|_{h^{2}}}{\|H\|_{h^{2}}}$ with respect to the transfer functions $H$ and $\hat{H}$ of the original system (35) and ROM, respectively. It should be noted that, while data-driven model reduction methods such as the Loewner framework and Data-driven Balanced Truncation (BT) are well-known, these methods require frequency response data or impulse response data as input. Therefore, they are challenging to apply directly in this context.
VI-B Numerical results
We present the results of applying Algorithm1 to the given state and input data. The initial point for Algorithm 1 is set to the reduced matrices obtained via DMDc, and $X^{(h_{x})},U^{(h_{u})},X^{(h_{x})}_{\mathrm{p}},\hat{A}$ satisfy Assumptions (A1)-(A6).
Figure 1 shows the $h^{2}$-norm plotted for different values of $r$ when the noise coefficient $\beta$ is 00. Similar to conventional model reduction methods, it is confirmed that larger values of $r$ lead to ROMs that better approximate the original system (35). Nevertheless, even at $r=400$ the relative $h^{2}$-error plateaus at about $15\%$ because an intrinsically nonlinear model is being approximated by a purely linear ROM.
Figure 2 illustrates the $h^{2}$-norm for different noise coefficients $\beta$ when $r=200$. When data satisfying the assumptions is used, a better ROM in terms of the $h^{2}$-norm is obtained compared to the initial point provided by DMDc. In the case of higher noise ($\beta=0.01$), the noise impact cannot be ignored, causing the gradient to deviate significantly from the true gradients, which results in limited improvement. Conversely, when the noise level is small ($\beta=0.001$), we observe a 50–60% improvement, similar to the no-noise case ($\beta=0$).
Figure 1: Convergence behavior of (10) with $r$
Figure 2: Convergence behavior of (10) with $\beta$
VII Concluding remarks
In this study, we developed a data-driven model reduction method for linear discrete-time systems that constructs optimal ROMs in the $h^{2}$-norm sense. Through both theoretical and numerical analyses, we demonstrated that if the measurement data snapshots satisfy some assumptions, an $h^{2}$-optimal ROM can be obtained compared to the initial reduced model. Moreover, when modeling changes in neural activity of C. elegans, our approach was shown to outperform DMDc in terms of achieving an improved ROM under the $h^{2}$-norm sense.
Several future directions are outlined below. First, we aim to construct more computationally efficient algorithms for large-scale problems. Second, we plan to report detailed results for additional application domains. Preliminary experiments indicate that the proposed method also outperforms DMDc when modeling daily new infection counts in epidemiology [33]. Third, we will extend (10) to an optimization problem on a Riemannian manifold. As shown in [12] and [44], the reduced matrix $\hat{A}$ obtained at each iteration is inherently stable, so the explicit stability constraint becomes unnecessary.
Acknowledgment
This work was supported by JSPS KAKENHI under Grant Numbers 23K03899 and 25KJ0986.
Proof of Lemma 1.
Since Assumption (A5), $X^{\dagger}X=I_{n}$. From [39], under Assumptions (A1) and (A3), (27) has a unique solution $R_{\text{data}}=R$. Similarly, under Assumptions (A2) and (A3), (28) has a unique solution $S_{\text{data}}=S$. $\Box$
Proof of Lemma 2.
Since (21) and (22) are solutions of (20), $\operatorname{rank}\bigl{(}\begin{smallmatrix}X&U\end{smallmatrix}\bigl{)}=\operatorname{rank}\bigl{(}\begin{smallmatrix}X&U&X_{\mathrm{p}}X^{\top}\end{smallmatrix}\bigl{)}$ holds. Thus, Assumption (A4) implies that (20) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$. $\Box$
Proof of Lemma 3.
(23) has the unique solution (24) when Assumption (A5) holds. From (32), we obtain $US_{B}=U_{B}S_{\text{data}}=UB^{\top}S_{\text{data}}$. Therefore, from Assumption (A6), $U^{\dagger}U=I_{m}$ holds. Thus, (33) holds. $\Box$
Proof of Theorem 1.
Since $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$ hold from Lemmas 2 and 3, the unique solutions $R$ and $S$ of (8) and (9), respectively, coincide with the unique solutions $R_{\text{data}}$ and $S_{\text{data}}$ of (27) and (28), as stated in Lemma 1.
Furthermore, from Lemma 3 and (8), we have $S^{\top}AR=S^{\top}(R-B\hat{B}^{\top})(\hat{A}^{-1})^{\top}=(S^{\top}R-S_{B}^{\top}\hat{B}^{\top})(\hat{A}^{-1})^{\top}$.
Therefore, under Assumptions (A1)-(A6), the gradients of Proposition 2 are rewritten as (29)-(31). $\Box$
Proof of Theorem 2.
To solve (20) for $(Z_{A},Z_{B})$, it is necessary to solve $NL-N$ linear equations of $(NL-N)\times(n+m)$. Considering the least-squares problem, the complexity is $\mathcal{O}((NL-N)^{2}(n+m))$. Similarly, the complexity required to solve (23) for $U_{B}$ is $\mathcal{O}((NL-N)^{2}n)$. The discrete‐time Lyapunov equations (6) and (7) are each solvable in $\mathcal{O}(r^{3})$ via the Bartels-Stewart method [39]. For the Sylvester equations (27) and (28), forming $X^{\dagger}$ and the matrix products $X^{\dagger}Z_{A}$, $X^{\dagger}Z_{B}\hat{B}^{\top}$, and $X^{\dagger}(X_{\mathrm{p}}-U_{B})$ each require $\mathcal{O}((NL-N)n^{2})$. With $X^{\dagger}X=I_{n}$, the resulting $n\times n$ Sylvester equations are solved in $\mathcal{O}(n^{3})$ using the Bartels-Stewart method [39], giving a total cost $\mathcal{O}((NL-N)n^{2}+n^{3})$. $\Box$